An object is projected upwards with a velocity of 33m/s from a point on the ground. What is the maximum height it reaches? What is the velocity at 50m above the ground? What is the time taken to reach 50m above ground? What is the total time of flight?


1
Expert's answer
2021-08-11T07:28:27-0400
"H=\frac{v^2}{2g}\\H=\frac{33^2}{2(9.8)}=55.6\ m\\
v^2=33^2-2(9.8)50\\v=10.4\frac{m}{s}"

"50=33t-(0.5)(9.8)t^2\\t=2.3\ s\\T=2\frac{33}{9.8}=6.7\ s"